# Knowledge in Superposition: Unveiling the Failures of Lifelong Knowledge Editing for Large Language Models

### v1

@@ Introduction
Further analysis of the interference term reveals a close relationship with superposition: if superposition does not exist in language models, the superposition term (Figure 1d) is zero, causing the interference term to vanish (Figure 1e), allowing for lossless lifelong editing. Superposition (Elhage et al. 2022b) refers to the situation where neural networks attempt to represent more features than the available dimensions. For example, if a simple neural network with three neurons (corresponding to a three-dimensional space) attempts to represent three features, each neuron can be assigned to one feature, forming an orthogonal basis in three-dimensional space (Figure 2a). But if the same network tries to represent six features (or more features), it will use superposition strategy, noisily encoding these features, where the directions corresponding to each feature will not be fully orthogonal (Figure 2b). Specifically, we find that the magnitude of interference term depends on the orthogonality of knowledge representations (Figure 1e). If these knowledge representations are perfectly orthogonal (non-superposition), the interference term is zero, enabling perfectly lossless knowledge editing, where only the target knowledge is updated without affecting unrelated knowledge. However, if they are not orthogonal (superposition), the interference will accumulate linearly and eventually tend toward infinity, leading to language models’ failure. In other words, whether superposition exists in language models is equivalent to whether lossless knowledge editing can be achieved, which will determine whether lifelong editing can be achieved.

### v2

@@ Introduction
Further analysis of the interference term reveals a close relationship with superposition: if superposition does not exist in language models, the superposition term (Figure 1d) is zero, causing the interference term to vanish (Figure 1e), allowing for lossless lifelong editing. Superposition (Elhage et al. 2022b) refers to the situation where neural networks attempt to represent more features than the available dimensions. For example, if a simple neural network with three neurons (corresponding to a three-dimensional space) attempts to represent three features, each neuron can be assigned to one feature, forming an orthogonal basis in three-dimensional space (Figure 2a). But if the same network tries to represent six features (or more features), it will use superposition strategy, noisily encoding these features, where the directions corresponding to each feature will not be fully orthogonal (Figure 2b). Specifically, we find that the magnitude of interference term depends on the orthogonality of knowledge representations (Figure 1e). If these knowledge representations are perfectly orthogonal (non-superposition), the interference term is zero, enabling perfectly lossless knowledge editing, where only the target knowledge is updated without affecting unrelated knowledge. However, if they are not orthogonal (superposition), the interference will accumulate linearly and eventually tend toward infinity, leading to language models’ failure (Appendix H). In other words, whether superposition exists in language models is equivalent to whether lossless knowledge editing can be achieved, which will determine whether lifelong editing can be achieved.
@@ Conclusion
Acknowledgements
This work is supported by the National Natural Science Foundation of China (No.U24A20335). This work is also supported by Beijing Natural Science Foundation (L243006) and the National Natural Science Foundation of China (No.62176257, No.62406321). This work is also supported by the Youth Innovation Promotion Association CAS and the China Postdoctoral Science Foundation under Grant Number 2024M753500.
@@ Acknowledgements
Appendix H H. Continuous Editing Causes Models’ Failure
In this section, we conduct experiments to illustrate the effects of lifelong editing based on the theoretical framework established in Section Expanding to Lifelong Editing. Specifically, we show that the continual introduction of interference to knowledge unrelated to the edited knowledge (the mathematical formulation of such interference is detailed in Section Expanding to Lifelong Editing) leads to the progressive forgetting of unrelated knowledge by the model.
In our experiments, we use ROME to edit the GPT-J. The dataset is divided into editing data and test data, where the editing data is used to continually modify the model, and the test data is used to record the confidence of unrelated knowledge throughout this process. The editing process continues until the model forgets unrelated knowledge associated with the test data. The experimental results are shown in Figure LABEL:fig:knowledge_forgetting_gpt-j.
From the results, it is evident that continual editing introduces persistent interference to unrelated knowledge, which progressively reduces the confidence of correct knowledge. As the editing process continues, this interference eventually results in the forgetting of unrelated knowledge.